# RoHOI: Robustness Benchmark for Human-Object Interaction Detection

### v1

@@ Abstract
Human-Object Interaction (HOI) detection is crucial for robot-human assistance, enabling context-aware support. However, models trained on clean datasets degrade in real-world conditions due to unforeseen corruptions, leading to inaccurate prediction. To address this, we introduce the first robustness benchmark for HOI detection, evaluating model resilience under diverse challenges. Despite advances, current models struggle with environmental variability, occlusion, and noise. Our benchmark, RoHOI, includes 20 corruption types based on HICO-DET and V-COCO datasets and a new robustness-focused metric. We systematically analyze existing models in the related field, revealing significant performance drops under corruptions. To improve robustness, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) strategy to guide the model to be optimized based on holistic and partial cues, dynamically adjusting the model’s optimization to enhance robust feature learning. Extensive experiments show our approach outperforms state-of-the-art methods, setting a new standard for robust HOI detection. Benchmarks, datasets, and code will be made publicly available at https://github.com/Kratos-Wen/RoHOI.
1 Introduction
Human-Object Interaction (HOI) detection is a key task in visual scene understanding, involving the identification and interpretation of interactions between humans and objects in an image [25, 19, 78, 56, 26, 20]. It serves as a foundation for many critical applications, including autonomous driving [45, 48], robotics [65], video surveillance [10, 50], and augmented reality [76], where a fine-grained understanding of human behavior and object interactions is paramount [13].
@@ 1 Introduction
Figure 1: Our RoHOI dataset comprises 20 types of algorithmically generated corruptions, systematically categorized into four groups, i.e., Optical system-induced artifacts, Sensor, compression and transmission artifacts, Environmentally induced artifacts, and Geometric and scene distortions. Each corruption type has five levels of severity, leading to a total of 100 distinct corruptions. These corruptions simulate real-world degradation factors encountered in practical applications.
In recent years, transformer-based models [78, 57, 82] and Vision-Language Models (VLMs) [2, 12, 8], have achieved impressive success in HOI detection field. These models excel at capturing intricate relationships between human-object pairs by leveraging large-scale datasets and contextual information. Despite these advancements, such models remain highly sensitive to various real-world corruptions, e.g., occlusion, lighting variations, and environmental shifts, which hinder their robustness and limit practical deployment. For instance, minor corruptions can cause substantial drops in accuracy, which poses significant risks in safety-critical applications like autonomous driving or robotics. This sensitivity raises serious concerns about the robustness and reliability of these models in unpredictable environments. Ensuring robust and reliable HOI detection in diverse, complex corruptions remains an unsolved challenge.
Existing HOI detection methods [78, 57, 2] predominantly evaluate model performance in idealized, clean conditions, overlooking real-world robustness.
This gap in the evaluation pipeline highlights the urgent need for a dedicated robustness benchmark. Such a benchmark can systematically evaluate model performance under diverse corruptions and perturbations, offering key insights into resilience and adaptability.
@@
In this paper, we introduce Robustness benchmark for HOI detection (RoHOI), systematically assessing HOI detection models under $20$ diverse corruptions designed to simulate real-world conditions based on HICO-DET [3] and V-COCO [13] datasets and robustness-focused metrics including Mean Robustness Index (MRI) [18] and our novel Composite Robustness Index (CRI). RoHOI extends robustness evaluation beyond clean datasets, exposing model vulnerabilities and guiding future improvements. Through our evaluations of existing HOI detection methods on the newly proposed robustness test sets, we observe that their performance remains limited. This finding underscores the necessity of developing more robust HOI detection models to ensure greater reliability in real-world applications.
@@
To mitigate the issue of performance degradation, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) method, which effectively leverages both holistic and partial cues from the clean training set. Specifically, we utilize SAM [28] to identify HOI-relevant regions and subsequently apply a dynamic masking strategy with varying ratios. This approach enables the model to focus on both holistic and partial information, thereby enhancing its robustness throughout the optimization process. Our proposed learning strategy delivers superior performance in HOI detection, achieving state-of-the-art (SOTA) performance on clean test set while delivering substantial improvements in recognition accuracy under the evaluation with the proposed corruptions. This advancement highlights the model’s enhanced resilience and reliability in challenging real-world conditions for HOI detection. The main contributions of this paper are summarized as follows:
@@
  • We for the first time introduce a robustness benchmark specifically tailored for HOI detection. Our benchmark incorporates $20$ distinct corruptions that closely mirror real-world challenges and a set of robustness-focused metrics.
@@
  • SAMPL yields SOTA in terms of robustness, surpassing prior HOI models under both clean and corrupted conditions. Our benchmark establishes a comprehensive evaluation framework for robustness in HOI detection.
2 Related Work
2.1 Human Object Interaction Detection
@@
Human-object interaction (HOI) detection [67, 11, 14, 32, 59, 47, 65, 69, 77, 81] seeks to automatically identify and interpret relationships between humans and objects in images. Existing approaches can be clustered into two major groups, which are one-stage and two-stage approaches.
Two-Stage HOI Detection. Two-stage methods [51, 64, 34, 80] rely on the detection results from the existing object detectors, e.g., Faster R-CNN [52]. CNN-based HOI detection methods [64, 80, 51] are explored by the researchers in the early stage. Graph structures [80, 64, 58, 61, 73] are well explored in the past to construct knowledge relation reasoning to pursue better human object interaction detection. Wan et al. [59] proposed pose aware multi-level network for HOI task. Zhang et al. [74] proposed an unpaired transformer architecture to achieve a better capture of the human-object interaction relations. Li et al. [33] proposed integrating and decomposing networks for HOI detection.
One-stage HOI Detection. One-stage HOI detection approaches [70, 75] have gained traction due to their efficiency and end-to-end learning advantages over two-stage methods. Liao et al. [35] introduced PPDM, leveraging parallel point detection and matching for real-time HOI detection. Kim et al. [24] proposed UnionDet, introducing a union-level detector to directly predict interaction triplets. Zhong et al. [79] introduced a glance-and-gaze mechanism to adaptively model action-aware points and refine interaction predictions. Recent advancements in transformer-based architectures further improved one-stage HOI detection. Kim et al. [27] proposed MUREN, which incorporates relational context learning to refine interaction reasoning. Yuan et al. [70] introduced RLIPv2, integrating Asymmetric Language-Image Fusion (ALIF) for scalable relational pretraining, achieving SOTA performance. VLMs have also been explored to enhance HOI detection, improving semantic understanding and generalization across unseen interactions. Liu et al. [23] proposed VLM-HOI, leveraging large-scale vision-language models for improved generalization. Zhang et al. [30] introduced EZ-HOI, enhancing zero-shot HOI triplet prediction through VLM-based feature learning. Addressing robustness in HOI detection, Wang et al. [7] proposed UAHOI, integrating uncertainty estimation to improve stability under varying conditions.
Most existing works are evaluated on clean datasets without considering data corruptions. However, in real-world applications, HOI detection models encounter various challenges leading to significant performance degradation. To address this, we introduce the first robustness benchmark for HOI detection, incorporating $20$ test-time corruptions.
2.2 Visual Robustness Benchmarks
Robustness is crucial for visual understanding as it ensures models can maintain high performance despite variations in environmental conditions, such as noise, occlusions, or lighting changes, which are common in real-world scenarios [4, 60, 53, 54, 1, 39, 68, 22, 44, 15, 63, 16, 40, 46, 71]. Researchers have constructed diverse robustness benchmarks to assess the capability of models coming from different visual understanding fields under diverse perturbations. Liu et al. [38] proposed ARES-Bench benchmark to evaluate the robustness of existing image classification methods under perturbations. Ishma et al. [21] constructed a large-scale benchmark for the visual question answer task. Hendrycks et al. [18] proposed ImageNet-C, a widely used benchmark assessing model robustness against common corruptions like noise, blur, and weather effects. Li et al. [31] proposed ImageNet-E to benchmark the model robustness in attribute editing field. Peng et al. [49] proposed new benchmark to asses robustness of existing one-shot skeleton based human action recognition models against occlusion perturbation. Chen et al. [9] constructed RoDLA to enable the evaluation of robust document layout analysis. However, none of the existing works focus on the robust evaluation of HOI detection methods. In this work, we open the vistas of the uncharted robust HOI detection area by constructing new RoHOI benchmark on two public HOI detection datasets and put forward the SAMPL method to achieve more reliable HOI detection performances.
3 Benchmark
3.1 Corruptions
To rigorously assess HOI detection robustness, we introduce our dedicated RoHOI benchmark, designed to evaluate model performance under structured real-world corruptions. While many corruptions in existing datasets [18], such as brightness and contrast shifts, are mitigated through standard augmentation, their weather-based effects (e.g., synthetic snow and frost) often fail to capture real-world optical complexities. Our benchmark introduces 20 corruption types, categorized into four groups: Optical system-induced (OS) artifacts, Sensor, compression, and transmission (SCT) artifacts , Environmentally induced (EI) artifacts, and Geometric and scene (G&S) distortions (see Fig. 1). Each corruption is applied at five severity levels, enabling a fine-grained analysis of robustness under increasing degradation.
@@ 3.1 Corruptions
OS Artifacts degrade spatial clarity and obscure interaction details by distorting object boundaries. Our benchmark includes motion blur (MB), defocus blur (DB), and Gaussian blur (GauB), which simulate degradation caused by camera movement, misfocus, and signal processing. We refine glass blur (GB) to better replicate distortions caused by imaging through uneven or scratched glass surfaces.
@@
SCT Artifacts stem from hardware limitations, lossy encoding, and data transmission errors. Alongside standard corruptions like Gaussian noise (GauN), shot noise (ShN), salt-and-pepper noise (S&P), JPEG artifacts (JPEG), and speckle noise (SN), we introduce packet loss (PL), which simulates data degradation due to incomplete transmissions.
@@
EI Artifacts alter image quality due to external conditions. Extreme lighting variations cause exposure (EXP) distortion. The rainbow effect (RE) misaligns color channels, distorting object hues. Occlusion (OCC) removes critical visual information when foreign objects block key scene elements, while vignette effect (VE) creates uneven brightness, reducing visibility at image borders.
@@
G&S Distortions alter spatial relationships and object proportions. Moire patterns (MP) create artificial wave-like distortions in high-frequency textures. Screen crack (SC) introduce irregular fractures and glare artifacts, while elastic transformation (ET) warp object structures. Perspective distortion (PD) skews object shapes when captured at extreme angles. Pixelation (PIX) replaces fine details with block-like artifacts due to low-resolution rendering, and zoom blur (ZB) degrades object sharpness through rapid motion along the optical axis.
3.2 Datasets
@@
To evaluate robustness under real-world corruptions, we use two widely adopted HOI detection datasets: HICO-DET and V-COCO. HICO-DET [62] contains $47,776$ images with $600$ interaction classes derived from $117$ action classes and $80$ object categories. Each image is annotated with human-object pairs and their corresponding interactions, making it a benchmark for evaluating diverse HOI scenarios. V-COCO [13] is a subset of MS-COCO [37] with $10,346$ images, $16,199$ human instances, $81$ object categories and $29$ verb classes. Unlike HICO-DET, which provides a fine-grained benchmark, V-COCO evaluates model performance on common human-object interactions. Evaluating these datasets under corruption provides a systematic robustness assessment beyond clean-image performance.
3.3 Baselines
@@
To systematically evaluate the robustness of HOI detection models, we benchmark both two-stage and one-stage architectures, including transformer-based and vision-language pretraining approaches. Two-stage methods first detect objects before predicting interactions. We evaluate QPIC [56] and UPT [74], both of which use object detectors to extract human-object proposals before interaction classification. We include CDN [72], which refines interaction reasoning via a cascade structure, and DiffHOI [66], which leverages text-to-image diffusion priors to improve rare interaction recognition. One-stage methods directly predict interactions without explicit object detection, improving efficiency. We evaluate FGAHOI [43], which enhances spatial alignment using fine-grained anchors, and QAHOI [5], which integrates query-based attention for interaction localization. Additionally, we include MUREN [27], which refines relational context modeling, and SOV-STG [6], which optimizes decoder attention for verb-object association. GEN-VLKT [36] incorporates visual-language pretraining to improve interaction recognition. RLIPv2 [70] incorporates region-language alignment for fine-grained interaction reasoning. Additionally, we develop SAMPL, which is built upon RLIPv2 and incorporates semantic-aware masking-based progressive learning. Our benchmark systematically evaluates these models across diverse corruption types and severity levels, providing a comprehensive assessment of HOI detection robustness.
3.4 Evaluation Metrics
@@
We evaluate clean test set of HICO-DET [62] using mAP at Intersection over Union (IoU) $0.5$ for all (Full), rare ($<$10 instances), and non-rare ($\geq$10 instances) categories [70]. For clean test set of V-COCO [13], we use $\mathrm{AP}_{\text{role}}$ for interaction detection, with $\mathrm{AP}^{\#1}_{\text{role}}$ for action recognition and $\mathrm{AP}^{\#2}_{\text{role}}$ requiring object localization at IoU $0.5$. To evaluate HOI detection performance on our robustness benchmark, we introduce two robustness-focused metrics: Mean Robustness Index (MRI) and Composite Robustness Index (CRI). MRI, following the definition in prior work [18], measures average performance under corruptions but overlooks stability across severity levels. To address this, we propose CRI, which penalizes performance variance while normalizing against clean-set accuracy, ensuring models with consistent robustness are appropriately rewarded.
3.4.1 Mean Robustness Index (MRI)
@@
MRI serves as a baseline measure of a model’s average performance across corruption types and severities:
@@
3.4.2 Composite Robustness Index (CRI)
To address MRI’s limitations, we propose CRI, which penalizes performance instability across corruption levels while normalizing against clean-set performance:
@@ 3.4.2 Composite Robustness Index (CRI)
where $\overline{mAP}c$ is the mean mAP for corruption type $c$, and $\sigma_{c}$ is the standard deviation across corruption levels. The clean-set performance $mAP{clean}$ is defined as mAP (Full) for HICO-DET and $\text{AP}^{\#2}_{\text{role}}$ for V-COCO, where $\mathrm{AP}^{\#2}_{\text{role}}$ requires both action and object localization, with an IoU threshold of 0.5. The penalty term $\frac{1}{\log(1+\sigma_{c})+1}$ reduces scores for models with high variance, ensuring that robustness accounts for both overall performance and stability.
@@
4 Method
In this section, we present our novel approach, Semantic-Aware Masking-based Progressive Learning (SAMPL), which improves model optimization by enhancing both holistic and partial feature learning. This is achieved through dynamically masked inputs, with the mask generation guided by the SAM [28] to identify Regions Of Interest (ROI) in HOI detection. The adaptive progressive learning strategy adjusts the masking procedure automatically based on the model’s performance on the validation set and the recorded frequency, ensuring more robust feature learning. We will now provide a detailed explanation of the design of our approach.
SAM Guided HOI-Aware Semantic Masking (SAMSM). Our goal is to maximize the utilization of the information provided by the clean training set to enhance the model’s robustness. This is achieved by ensuring that both holistic and partial cues are effectively extracted and learned, ultimately benefiting model optimization. By applying masks of varying scales to the semantic regions of interest in the image (i.e., areas related to interacted objects and person areas), we assist the model in learning HOI cues within the restricted regions.
To implement this masking strategy, we use ground-truth bounding boxes and SAM-generated masks for instance-aware occlusions. Each object’s bounding box is processed by SAM to obtain a fine-grained segmentation mask $\mathbf{M}_{b}$. To prevent reliance on instance contours, we apply random edge expansion, ensuring occlusion does not strictly follow the instance’s boundaries. Instead of naively occluding entire bounding boxes, we selectively mask regions within $\mathbf{M}_{b}$ while preserving surrounding contextual cues. This strategy prevents the model from relying solely on object appearance and encourages it to infer interactions based on partial information. During training, the masked pixels in the image are replaced with zero values.
Next, we define various severity levels for the mask to obtain different partial cues from the original image. These severity levels are represented by $\Omega=\{\omega_{1},\omega_{2},\omega_{3},\omega_{4}\}$, corresponding to clean, low, middle, and high levels. The clean level represents the original, unmasked data to support holistic feature learning. For the other three levels, cover ratios are determined to control the final height and width of the generated black box for semantic-aware masking, which can be denoted as $\left[r^{w}_{\omega},r^{h}_{\omega}\right]$, where $\omega\in\left[\omega_{2},\omega_{3},\omega_{4}\right]$.
We then generate the mask for the severity level $\omega$ using the following equation:
@@ 4 Method
In this process, we first apply binary dilation to the semantic mask $\mathbf{M}_{b}$, then compute the rectangular convex hull, and finally resize the resulting rectangular mask using the corresponding ratio $\left[r^{w}_{\omega},r^{h}_{\omega}\right]$. The operation $\Phi$ refers to the resizing procedure that controls the mask size. The final mask, $\mathbf{M}^{b}_{\omega}$, is the resulting mask that aids partial feature learning in our proposed robust training method.
@@
FGAHOI [43] Swin-T 29.81 22.17 32.09 16.05 0.22 - - - -
GEN-VLKT-S [36] R50 33.75 29.25 35.10 17.01 0.22 62.41 64.46 34.69 0.19
GEN-VLKT-L [36] R101 34.95 31.18 36.08 18.93 0.23 63.58 65.93 37.45 0.20
QAHOI [5] Swin-T 28.47 22.44 30.27 15.07 0.25 58.23 58.70 29.33 0.20
SOV-STG-S [6] R50 33.80 29.28 35.15 16.94 0.22 62.88 64.47 34.40 0.20
MUREN [27] R50 32.88 28.70 34.13 16.32 0.23 68.10 69.88 37.10 0.18
RLIPv2-ParSeDA [70] Swin-T 38.60 33.66 40.07 24.58 0.28 68.83 70.76 45.05 0.24
SAMPL (ours) Swin-T 39.16 33.89 40.73 24.25 0.29 69.04 71.13 48.83 0.27

### v2

@@ Abstract
Human-Object Interaction (HOI) detection is crucial for robot-human assistance, enabling context-aware support. However, models trained on clean datasets degrade in real-world conditions due to unforeseen corruptions, leading to inaccurate prediction. To address this, we introduce the first robustness benchmark for HOI detection, evaluating model resilience under diverse challenges. Despite advances, current models struggle with environmental variability, occlusions, and noise. Our benchmark, RoHOI, includes 20 corruption types based on the HICO-DET and V-COCO datasets and a new robustness-focused metric. We systematically analyze existing models in the HOI field, revealing significant performance drops under corruptions. To improve robustness, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) strategy to guide the model to be optimized based on holistic and partial cues, thus dynamically adjusting the model’s optimization to enhance robust feature learning. Extensive experiments show that our approach outperforms state-of-the-art methods, setting a new standard for robust HOI detection. Benchmarks, datasets, and code will be made publicly available at https://github.com/Kratos-Wen/RoHOI.
 [^1] The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation). This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026).
@@
Figure 1: Our RoHOI dataset comprises 20 types of algorithmically generated corruptions, systematically categorized into four groups. Each corruption type has five levels of severity, leading to a total of 100 distinct corruptions. These corruptions simulate realistic semantic and structural disturbances specifically encountered in practical application.
Introduction
Human-Object Interaction (HOI) detection aims to identify and interpret interactions between humans and objects (Kim et al. 2021; Hou et al. 2021a; Zhang et al. 2022; Tamura, Ohashi, and Yoshinaga 2021; Kim et al. 2022; Hou et al. 2021b). It serves many critical applications, including autonomous driving (Martin et al. 2019; Peng et al. 2022), robotics (Xu et al. 2020), video surveillance (Dogariu et al. 2020; Prest, Ferrari, and Schmid 2013), and augmented reality (Zhang, Cui, and Fang 2025), where a fine-grained understanding of human behavior and object handling is paramount (Gupta and Malik 2015).
In recent years, transformer-based models (Zhang et al. 2022; Tu et al. 2023; Zou et al. 2021) and Vision-Language Models (VLMs) (Cao et al. 2023; Gao et al. 2024; Chen et al. 2024a), have achieved impressive success in the HOI detection field. These models excel at capturing intricate relationships between human-object pairs by leveraging large-scale datasets and contextual information.
Despite these advancements, existing models remain sensitive to real-life corruptions, particularly semantic distortions like occlusions and perspective shifts, as well as environmental variations. Even minor disturbances may pose significant risks in critical applications like autonomous driving and robotics. Enhancing robustness under these conditions remains a crucial but unresolved challenge. Existing HOI detection methods (Zhang et al. 2022; Tu et al. 2023; Cao et al. 2023) had their performance evaluated predominantly under ideal conditions. Thus, a dedicated robustness benchmark tailored explicitly to HOI detection, covering both pixel-level and semantic-level corruptions, is urgently required to put their resilience to the test.
@@ Introduction
In this paper, we introduce the Robustness benchmark for HOI detection (RoHOI), systematically assessing HOI detection models under $20$ diverse forms of data corruption. These were designed to simulate real-world conditions. The HICO-DET (Chao et al. 2018) and V-COCO (Gupta and Malik 2015) datasets served as starting point. We include robustness-focused metrics, such as the Mean Robustness Index (MRI) (Hendrycks and Dietterich 2019) and our novel Composite Robustness Index (CRI). Through our evaluations of existing HOI detection methods on the newly proposed robustness test set, we observe that their performance is less than optimal. This finding underscores the need to develop more robust HOI detection models.
@@
To mitigate the issue of performance degradation, we propose our Semantic-Aware Masking-based Progressive Learning (SAMPL) method. Specifically, we utilize SAM (Kirillov et al. 2023) to identify HOI-relevant regions and subsequently apply a dynamic masking strategy with varying size ratios. This approach enables the model to focus on both holistic and partial information, thereby enhancing its robustness throughout the optimization process.
Our proposed learning strategy delivers superior HOI detection, retaining state-of-the-art (SOTA) performance on a clean test set while delivering substantial improvements under corruptions. This highlights the model’s enhanced resilience and reliability in challenging real-world conditions.
The main contributions of this paper are summarized as follows:
@@
  • We for the first time introduce a robustness benchmark specifically tailored to HOI detection. Our benchmark incorporates $20$ distinct corruptions that closely mirror real-world challenges and a set of robustness-focused metrics.
@@
  • SAMPL yields SOTA in terms of robustness, surpassing prior HOI models under both clean and corrupted conditions.
Related Work
@@
Human Object Interaction Detection. Human-object interaction (HOI) detection (Yao and Fei-Fei 2010; Fang et al. 2018; Gupta, Schwing, and Hoiem 2019; Li et al. 2019; Wan et al. 2019; Park et al. 2018; Xu et al. 2020; Yu, Liu, and Odobez 2018; Zhang et al. 2015; Zhou et al. 2020) aims to automatically identify and interpret relationships between humans and objects, typically divided into two categories: two-stage and one-stage approaches.
Two-stage HOI detection methods (Qi et al. 2018; Xu et al. 2019; Liang et al. 2021; Zhou and Chi 2019) first detect humans and objects independently before classifying interactions. Early CNN-based approaches (Xu et al. 2019; Zhou and Chi 2019; Qi et al. 2018) mainly utilized spatial contexts for interaction modeling. Later graph-based methods (Zhou and Chi 2019; Xu et al. 2019; Ulutan, Iftekhar, and Manjunath 2020; Wang et al. 2021; Zhang, Campbell, and Gould 2021) improved relational reasoning explicitly. Subsequent methods further enhanced precision via pose-aware representations (Wan et al. 2019), transformer-based unpaired modeling (Zhang, Campbell, and Gould 2022), and integration-decomposition strategies (Li et al. 2020).
One-stage HOI detection methods (Yuan et al. 2023; Zhang et al. 2023) achieve efficiency by directly predicting human-object interactions end-to-end. Early methods exploited parallel point detection (Liao et al. 2020) and union-level modeling (Kim et al. 2020), while subsequent approaches refined interaction localization via adaptive glance-and-gaze mechanisms (Zhong et al. 2021). Transformer-based methods enhanced relational context modeling (Kim, Jung, and Cho 2023) and language-image fusion (RLIPv2)(Yuan et al. 2023), significantly improving performance. Recently, vision-language models (VLMs) were leveraged to boost generalization across unseen interactions (VLM-HOI(Kang et al. 2024)) and zero-shot triplet predictions (EZ-HOI (Lei, Wang, and Tan 2024)). Robustness-specific methods further integrated uncertainty estimation (UAHOI (Chen, Chen, and Yang 2024)) to improve stability under perturbations. Nevertheless, all these approaches primarily evaluate performance on clean data, neglecting real-world corruptions. To address this gap, we introduce RoHOI, the first HOI robustness benchmark encompassing $20$ test-time corruptions.
Visual Robustness Benchmarks. Robustness benchmarks have been widely established to assess visual model resilience across various tasks (Chattopadhyay et al. 2021; Wang et al. 2024; Sarkar et al. 2024; Schiappa et al. 2024; An et al. 2024; Liu et al. 2024b; Yi et al. 2024; Kamann and Rother 2020; Ma et al. 2024b; Hao et al. 2024; Xie et al. 2025; Hao et al. 2025; Liu et al. 2024c; Michaelis et al. 2019; Zeng et al. 2024), including image classification (ARES-Bench (Liu et al. 2024a)), visual question answering (Ishmam et al. 2024), common corruption evaluation (ImageNet-C (Hendrycks and Dietterich 2019)), attribute editing robustness (ImageNet-E (Li et al. 2023)), one-shot skeleton-based action recognition under occlusion (Peng et al. 2023), and robust document layout analysis (RoDLA (Chen et al. 2024b)). However, robustness benchmarks explicitly designed for HOI detection have not been explored. In this work, we address the uncharted robust HOI detection area by constructing the new RoHOI benchmark and by putting forward the SAMPL method to achieve more reliable HOI detection performance.
Benchmark
Corruptions. To rigorously assess the robustness of HOI detection, we introduce our dedicated RoHOI benchmark, designed to evaluate model performance under real-world corruptions. Existing benchmarks such as ImageNet-C (Hendrycks and Dietterich 2019) primarily cover low-level pixel disturbances like brightness or blur. In contrast, RoHOI explicitly includes semantic-level corruptions specifically tailored to challenge spatial and relational reasoning critical for HOI detection. RoHOI explicitly introduces 20 corruptions derived from HICO-DET and V-COCO images, categorized into four groups: optical system artifacts such as motion blur and glass blur; sensor, compression, and transmission distortions such as packet loss and JPEG artifacts; environmental effects including exposure imbalance and occlusion; and geometric or structural disruptions such as screen cracks, elastic deformation, and perspective distortion. The corruptions range from pixel-level noise to semantic-level challenges that affect spatial alignment and interaction reasoning. Each corruption is implemented with five severity levels to support detailed analysis under progressively more difficult conditions.
@@
Optical System-induced (OS) Artifacts degrade spatial clarity and obscure interaction details by distorting object boundaries. Our benchmark includes motion blur (MB), defocus blur (DB), and Gaussian blur (GauB), which simulate degradation caused by camera movement, defocus, and signal processing. We refine glass blur (GB) to better replicate distortions caused by imaging through uneven or scratched glass surfaces.
@@
Sensor, Compression and Transmission (SCT) Artifacts stem from hardware limitations, lossy encoding, and data transmission errors. Alongside standard corruptions like Gaussian noise (GauN), shot noise (ShN), salt-and-pepper noise (S&P), JPEG artifacts (JPEG), and speckle noise (SN), we introduce packet loss (PL), which simulates data degradation due to incomplete transmissions.
@@
Environmentally Induced (EI) Artifacts alter image quality due to external conditions. Extreme lighting variations cause exposure (EXP) distortion. The rainbow effect (RE) misaligns color channels, distorting object hues. Occlusion (OCC) removes critical visual information when foreign objects block key scene elements, while the vignette effect (VE) creates uneven brightness, reducing visibility at image borders.
@@
Geometric and Scene (G&S) Distortions alter spatial relationships and object proportions. Moire patterns (MP) create artificial wave-like distortions in high-frequency textures. Screen crack (SC) introduce irregular fractures and glare artifacts, while elastic transformation (ET) warp object structures. Perspective distortion (PD) skews object shapes when captured at extreme angles. Pixelation (PIX) replaces fine details with block-like artifacts due to low-resolution rendering, and zoom blur (ZB) degrades object sharpness through rapid motion along the optical axis.
@@
Datasets. To evaluate robustness under real-world corruptions, we use two widely adopted HOI detection datasets: HICO-DET and V-COCO. HICO-DET (Wang et al. 2020) contains $47,776$ images with $600$ interaction classes derived from $117$ action classes and $80$ object categories. Each image is annotated with human-object pairs and their corresponding interactions, making it a benchmark for evaluating diverse HOI scenarios. V-COCO (Gupta and Malik 2015) is a subset of MS-COCO (Lin et al. 2014) with $10,346$ images, $16,199$ human instances, $81$ object categories and $29$ verb classes. Unlike HICO-DET, which provides a fine-grained benchmark, V-COCO evaluates model performance on common human-object interactions. Evaluating these datasets under corruption provides a systematic robustness assessment beyond clean-image performance.
@@
Baselines. To systematically evaluate the robustness of HOI detection models, we benchmark both two-stage and one-stage architectures, including transformer-based and vision-language pretraining approaches. Two-stage methods first detect objects before predicting interactions. We evaluate QPIC (Tamura, Ohashi, and Yoshinaga 2021) and UPT (Zhang, Campbell, and Gould 2022), both of which use object detectors to extract human-object proposals before interaction classification. We include CDN (Zhang et al. 2021), which refines interaction reasoning via a cascade structure, and DiffHOI (Yang et al. 2023), which leverages text-to-image diffusion priors to improve rare interaction recognition. One-stage methods directly predict interactions without explicit object detection, improving efficiency. We evaluate FGAHOI (Ma et al. 2024a), which enhances spatial alignment using fine-grained anchors, and QAHOI (Chen and Yanai 2023), which integrates query-based attention for interaction localization. Additionally, we include MUREN (Kim, Jung, and Cho 2023), which refines relational context modeling, and SOV-STG (Chen, Wang, and Yanai 2023), which optimizes decoder attention for verb-object association. GEN-VLKT (Liao et al. 2022) incorporates visual-language pretraining to improve interaction recognition. RLIPv2 (Yuan et al. 2023) incorporates region-language alignment for fine-grained interaction reasoning. Additionally, we develop SAMPL, which is built upon RLIPv2 and incorporates semantic-aware masking-based progressive learning. Our benchmark systematically evaluates these models across diverse corruption types and severity levels, providing a comprehensive assessment of HOI detection robustness.
@@
Evaluation Metrics. We evaluate the clean test set of HICO-DET (Wang et al. 2020) using mAP at Intersection over Union (IoU) $0.5$ for all (Full), rare (<<10 instances), and non-rare ($\geq$10 instances) categories (Yuan et al. 2023). For the clean test set of V-COCO (Gupta and Malik 2015), we use $\mathrm{AP}_{\text{role}}$ for interaction detection, with $\mathrm{AP}^{\#1}_{\text{role}}$ for action recognition and $\mathrm{AP}^{\#2}_{\text{role}}$ requiring object localization at IoU $0.5$. To evaluate HOI detection performance on our robustness benchmark, we introduce two robustness-focused metrics: Mean Robustness Index (MRI) and Composite Robustness Index (CRI). MRI, following the definition in prior work (Hendrycks and Dietterich 2019), measures average performance under corruptions but overlooks stability across severity levels. To address this, we propose CRI, which penalizes performance variance while normalizing against clean-set accuracy, ensuring models with consistent robustness are appropriately rewarded.
@@
Mean Robustness Index (MRI) serves as a baseline measure of a model’s average performance across corruption types and severities:
@@
Composite Robustness Index (CRI) addresses MRI’s limitations. It penalizes performance instability across corruption levels while normalizing against clean-set performance:
@@
where $\overline{mAP}_{c}$ is the mean mAP for corruption type $c$, and $\sigma_{c}$ is the standard deviation across corruption levels. The clean-set performance $mAP_{clean}$ is defined as mAP (Full) for HICO-DET and $\mathrm{AP}^{\#2}_{\text{role}}$ for V-COCO, where $\mathrm{AP}^{\#2}_{\text{role}}$ requires both action and object localization, with an IoU threshold of 0.5. The penalty term $\frac{1}{\log(1+\sigma_{c})+1}$ reduces scores for models with high variance, ensuring that robustness accounts for both overall performance and stability.
@@
Method
In this section, we present our novel Semantic-Aware Masking-based Progressive Learning (SAMPL), aiming to enhance model robustness by progressively guiding models from holistic feature reliance towards robust contextual inference. Unlike standard augmentation methods, SAMPL employs adaptive semantic masking guided by SAM (Kirillov et al. 2023), dynamically adjusting masking severity based on validation performance. This structured semantic-level masking strategy explicitly promotes robust contextual reasoning by training models to infer interactions from progressively restricted semantic regions, thus effectively improving general robustness against a variety of realistic disturbances. Moreover, SAMPL’s structured progressive curriculum inherently regularizes feature representation, effectively reducing overfitting to dataset biases, thereby further improving generalization under unforeseen corruptions.
SAM Guided HOI-Aware Semantic Masking (SAMSM). To fully leverage clean training data, we generate instance-aware masks based on ground-truth bounding boxes and SAM-derived fine-grained masks ($\mathbf{M}_{b}$). Random edge expansions are applied to prevent strict reliance on instance contours, selectively masking semantic regions while preserving contextual cues. During training, masked pixels are set to zero, compelling the model to infer interactions from partial visual information.
We define various severity levels for the mask, represented by $\Omega=\{\omega_{1},\omega_{2},\omega_{3},\omega_{4}\}$, corresponding to clean, low, middle, and high levels. Clean level $\omega_{1}$ employs unmasked data for holistic feature learning, while other levels apply increasing occlusion severity via controlled cover ratios $\left[r^{w}_{\omega},r^{h}_{\omega}\right]$. Each semantic mask $\mathbf{M}^{b}_{\omega}$ is computed as:
@@
where dilation and convex hull operations refine the mask shape, and $\Phi$ denotes resizing. This structured masking approach not only simulates realistic semantic occlusions but also inherently promotes context-aware reasoning, significantly enhancing robustness against various semantic and structural corruptions.